OpenStudy (anonymous):

convert r=5csc theta to rectangular equation

OpenStudy (tkhunny):

Have you considered a slight transformation? \(r = 5\csc(\theta) = \dfrac{5}{\sin(\theta)} \rightarrow r\cdot\sin(\theta) = 5\)

OpenStudy (anonymous):

no i never thought about putting it that way

OpenStudy (tkhunny):

This is why you endlessly studied trig identities. It is time to prove it.